4. What magnitude magnetic field would be required to store ten kilowatt-hours (36 MJ) of magnetic potential energy in each cubic meter of vacuum or nonmagnetic material? ONE EQUATION USED SOLUTION ANSWER luse μ- and u = 5·The electric flux is (4.0 V·m/s5r5 through a certain area of a dielectric. Its dielectric constant is 2.5 so its permittivity is 2.2 x 10-11 F/m. Find the displacement current through that area at EQUATION USED 1.5 s. SOLUTION (SHOW ALL YOUR WORK.) ANSWER
Answer #1

4.

We know that, Energy stored per unit volume can be given as :

u = B2/2\mu0

Given: u = 36 MJ

Assuming , \mu0 = 4\pi×10-7 H

Putting all values ,

36×106 = B2/ (2×4×3.14×10-7)

By solving we get

B = 9.51 T